# vApps: Verifiable Applications at Internet Scale

### v2

@@ Abstract.
Blockchain technology promises a decentralized, trustless, and interoperable infrastructure. However, widespread adoption remains hindered by issues such as limited scalability, high transaction costs, and the complexity of maintaining coherent verification logic across different blockchain layers. This paper introduces Verifiable Applications (vApps), a novel development framework designed to streamline the creation and deployment of verifiable blockchain computing applications. vApps offer a unified Rust-based Domain-Specific Language (DSL) within a comprehensive SDK, featuring modular abstractions for verification, proof generation, and inter-chain connectivity. This eases the developer’s burden in securing diverse software components, allowing them to focus on application logic. The DSL also ensures that applications can automatically take advantage of specialized precompiles and hardware acceleration to achieve consistently high performance with minimal developer effort, as demonstrated by benchmark results for zero-knowledge virtual machines (zkVMs). Experiments show that native Rust execution eliminates interpretation overhead, delivering up to an 832x cycle count improvement compared to EVM-based approaches. Precompiled circuits can accelerate the proof by more than 95%, while GPU acceleration increases throughput by up to 30x and recursion compresses the proof size by up to 230x, enabling succinct and efficient verification. The framework also supports seamless integration with the Web2 and Web3 systems, enabling developers to focus solely on their application logic. Through modular architecture, robust security guarantees, and composability, vApps pave the way toward a trust-minimized and verifiable Internet-scale application environment.
@@ 5.2. SDK Examples
A minimal version of vApp in Rust DSL is presented here as an illustration for following the S0-S4 steps of a program’s transaction lifecycle as discussed in Section 4.2. This practice facilitates the proof generation as introduced in Section 6 to benefit from specialized precompiles and hardware acceleration. The correspondence with the transaction lifecycle is highlighted through green-colored comments in the code script, as illustrated in Listing 1 below.
@@ 6.3. Performance
Due to their highly modular nature, zkVMs are suitable for optimizations at the proof system level and for hardware acceleration. In this section, we describe techniques for such optimization. First, we map the distribution of cycle counts to steps S0-S4 of the transaction lifecycle (Section 4.2) for a program that proves Ethereum execution with Reth. We then show how these cycle counts can be optimized via specialized precompiles for various cryptographic operations, how the EVM can be repriced to account for ZK proving costs, and how hardware acceleration with GPUs can improve performance. We conduct these benchmarks on SP1 (succinct_sp1, ).
@@
We map the individual steps in proving Ethereum blocks to stages S0 to S4 from Figure 8. This mapping demonstrates that Ethereum execution follows the standard vApp transaction lifecycle. We see that the zkVM spends a large fraction of its cycle count (approaching upwards of in some cases), in the block execution phase, corresponding to S2, the state transition function. For small blocks, only 28.28% of cycles are spent on execution, as the cost of data access is not well amortized across a small number of transactions. We note that recover_senders is optimized via an ECDSA recovery precompile.
@@
deserialize_inputs (S0) | 13.26 % | 17.81 % | 24.07 % |
recover_senders (S0) | 1.97 % | 2.86 % | 5.50 % |
block_validation (S0 and S4) | 1.18 % | 1.41 % | 1.92 % |
initialize_witness_db (S1) | 13.48 % | 18.05 % | 24.08 % |
block_execution (S2 and S4) | 63.33 % | 50.47 % | 28.28 % |
state_root_computation (S3) | 5.82 % | 8.23 % | 14.57 % |
accrue_logs_bloom (S4) | 0.57 % | 0.64 % | 0.75 % |
other (S0 and S4) | 0.38 % | 0.53 % | 0.83 % |

### v3

@@ Abstract.
Blockchain technology promises a decentralized, trustless, and interoperable infrastructure. However, widespread adoption remains hindered by issues such as limited scalability, high transaction costs, and the complexity of maintaining coherent verification logic across different blockchain layers. This paper introduces Verifiable Applications (vApps), a novel development framework designed to streamline the creation and deployment of verifiable blockchain computing applications. vApps offer a unified Rust-based Domain-Specific Language (DSL) within a comprehensive SDK, featuring modular abstractions for verification, proof generation, and inter-chain connectivity. This eases the developer’s burden in securing diverse software components, allowing them to focus on application logic. The DSL also ensures that applications can automatically take advantage of specialized precompiles and hardware acceleration to achieve consistently high performance with minimal developer effort, as demonstrated by benchmark results for zero-knowledge virtual machines (zkVMs). Experiments show that native Rust execution eliminates interpretation overhead, delivering up to an 832x cycle count improvement compared to EVM-based approaches. Precompiled circuits can accelerate proving by more than 95%, while GPU acceleration increases throughput by up to 30x and recursion compresses the proof size by up to 230x, enabling succinct and efficient verification. The framework also supports seamless integration with Web2 and Web3 systems, enabling developers to focus solely on their application logic. Through modular architecture, robust security guarantees, and composability, vApps pave the way toward a trust-minimized and verifiable Internet-scale application environment.
@@ 5.2. SDK Examples
A minimal version of a vApp in the Rust DSL is presented here as an illustration for following the steps S0-S4 of a program’s transaction lifecycle as discussed in Section 4.2. This practice facilitates the proof generation as introduced in Section 6 to benefit from specialized precompiles and hardware acceleration. The correspondence with the transaction lifecycle is highlighted through green-colored comments in the code script, as illustrated in Listing 1 below.
@@ 6.3. Performance
Due to their highly modular nature, zkVMs are suitable for optimizations at the proof system level and for hardware acceleration. In this section, we describe techniques for such optimization. First, we map the distribution of cycle counts to steps S0-S4 of the transaction lifecycle (Section 4.2) for a program that proves Ethereum execution with Reth. We then show how these cycle counts can be optimized via specialized precompiles for various cryptographic operations, how the EVM can be repriced to account for ZK proving costs, how zkVMs can avoid interpreter overhead with native code execution, and how hardware acceleration with GPUs can improve performance. We conduct these benchmarks on SP1 (succinct_sp1, ).
@@
We map the individual steps in proving Ethereum blocks to stages S0 to S4 from Figure 8. This mapping demonstrates that Ethereum execution follows the standard vApp transaction lifecycle. We see that the zkVM spends a large fraction of its cycle count (approaching upwards of in some cases), in the block execution phase, corresponding to S2, the state transition function. For small blocks, only 28.96% of cycles are spent on execution, as the cost of data access is not well amortized across a small number of transactions. We note that recover_senders is optimized via an ECDSA recovery precompile.
@@
| deserialize_inputs (S0) | 12.66 % | 16.67 % | 22.47 % |
| recover_senders (S0) | 2.16 % | 3.08 % | 5.93 % |
| block_validation (S0 and S4) | 1.10 % | 1.27 % | 1.73 % |
| initialize_witness_db (S1) | 13.39 % | 17.63 % | 23.64 % |
| block_execution (S2 and S4) | 63.31 % | 51.29 % | 28.96 % |
| state_root_computation (S3) | 6.30 % | 8.72 % | 15.43 % |
| accrue_logs_bloom (S4) | 0.52 % | 0.58 % | 0.68 % |
| other (S0 and S4) | 0.56 % | 0.76 % | 1.15 % |